OpenStudy (anonymous):

what is the sum of the measures of the interior angles of a 14- sided polygon im thinking its 2,160 but cant explain it?

jigglypuff314 (jigglypuff314):

sum of measures of interior angles of n-sided polygon = (n-2)180

jigglypuff314 (jigglypuff314):

so ( (14) - 2) times 180 =

OpenStudy (anonymous):

thanks

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

2,160 right

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (anonymous):

yeyyyyyy

OpenStudy (anonymous):

im happy nowwww lol I love this site

jigglypuff314 (jigglypuff314):

lol I'm glad :)

OpenStudy (anonymous):

can u help me on another one please

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

ok lina is covering her

OpenStudy (anonymous):

attic wall with wallpaper the wall is a trapezoid-shaped with a top and bottom of 14 feet and 20 feet the wall is 8 feet how much more will she need to cover up her wall

OpenStudy (anonymous):

of bases*

jigglypuff314 (jigglypuff314):

area of trapezoid = average of bases times height

OpenStudy (anonymous):

I kind of guessed 163 but I cant do the problem I keep getting a decimal but my multiple choice doesn't have decimals in the answers

jigglypuff314 (jigglypuff314):

what is the average of 14 and 20? (14+20) divided by 2 =

OpenStudy (anonymous):

24

jigglypuff314 (jigglypuff314):

then times 8

jigglypuff314 (jigglypuff314):

no. wait... (14+20) = 34 34 / 2 = 17

OpenStudy (anonymous):

192

OpenStudy (anonymous):

ohhhh crap yeah sorry

OpenStudy (anonymous):

times 8 =136

jigglypuff314 (jigglypuff314):

yup :) does that work with your multiple choice?

OpenStudy (anonymous):

YUSHHHHHHH LMAOOO

OpenStudy (anonymous):

THANKKKK YOUUUU

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

wow I suck at math im gonna be posting up some more questions

jigglypuff314 (jigglypuff314):

practice makes perfect :)

OpenStudy (anonymous):

true but I don't want to bother anyone anymore I think I bothered enough

jigglypuff314 (jigglypuff314):

it's fine XD I've spent 10 hours straight on this site before, helping people out :P but ya know... there's this "Best Response" button over there --> It really encourages the helpers on this site to keep working for free xD ^

OpenStudy (anonymous):

cool lol cuz I really need help geometry is easy but we still didn't cover everything up lol

jigglypuff314 (jigglypuff314):

well then post your question and I'll see what I can do to help :)

OpenStudy (anonymous):

yeyyyy